OpenStudy (anonymous):

Totally lost: integrate(x/(x^2+y^2)dx)

OpenStudy (anonymous):

let u=x^2+y^2

OpenStudy (anonymous):

\[\int\limits_{}^{}x/(x^2+y^2)\] , then what is dv?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

Sorry, I'm still lost. I'm letting u=x^2+y^2, so du=2x, I got that part, but what do dv and v become?

OpenStudy (anonymous):

Ah! Got it, u-substitiutuion, not integration by parts! Thanks.

OpenStudy (anonymous):

\[\int\limits \frac{x}{x^2+y^2} \, dx=\frac{1}{2} \text{Log}\left[x^2+y^2\right]+c \]